Kim(x,y)->(0,0) sinx/y exist.is it true or false. Give reasons for your answer


1
Expert's answer
2021-02-03T15:23:36-0500

Solution:

"\mathrm{Let's \ find\:two\:different\:paths\:to\:approach\:the\:point\:that}"

"\mathrm {\:gives\:different\:values\:for\:the\:limit}"

Case 1: "\lim_{(x,y)\rightarrow (0,0)}(\sin (\dfrac{x}{y}))" along "x=0" :

"=\sin (\dfrac{0}{y})=\sin (0)=0"

Case 2: "\lim_{(x,y)\rightarrow (0,0)}(\sin (\dfrac{x}{y}))" along "x=y" :

"=\sin (\dfrac{y}{y})=\sin (1)"

Thus, from case 1 and 2, we have that given limit diverges (or does not exist).

Hence, the given statement is false.